Solve each of the following systems by using either addition or substitution. If a unique solution does not exist, state whether the system is dependent or inconsistent.
10x +2y = 7
Y = -5x + 3

Answer by funmath(2873)

Solve each of the following systems by using either addition or substitution. If a unique solution does not exist, state whether the system is dependent or inconsistent.
E1) 10x +2y = 7
E2) Y = -5x + 3
Let the y in E1=-5x+3:
10x+2(-5x+3)=7
10x-10x+6=7
6=7
The variable disappeared and 6 does not = 7, therefore there is no solution and the system is inconsistent.
Happy Calculating!!!